Question 2 (1 point)

Boulder Mountain Ski Company has total assets of $497,600,000 and a debt ratio of 0.26. Calculate the company’s debt-to-equity ratio. Round to two decimal places.

Question 3 (1 point)

The Timber Ridge Company has the following relationships:

Sales/Total assets = 2.90; ROA = 0.0920

What is Timber Ridge’s net profit margin? Round to 3 decimal places.

Question 5 (1 point)

Archware Systems has total assets of $35.594
billion, total debt of $9.678 billion, and net sales of $22.510 billion. Their
net profit margin for the year was 0.22, while the operating profit
margin was 30 percent. What is Archware’s net income? (Answer needs to be stated in billions. For example: 2.83) Round to two decimal places.

Question 7 (1 point)

Blue Air Inc., has net sales of $702,000 and accounts receivables of $167,000. What are the firm's accounts receivables turnover? Round to two decimal places.

Question 8 (1 point)

JP Vineyards has sales of $828,000, a gross profit margin of 0.322, and inventory of $177,000. What is the company's inventory turnover ratio? Round to two decimal places.
Answer #1

2

D/E =D/A/(1-D/A) = 0.26/(1-0.26) = 0.35

3

Profit margin = ROA/(sales/total assets) = 0.092/2.9=0.032

Please ask remaining parts seperately, questions are unrelated, I have done one bonus